OpenStudy (anonymous):

Use the graph below for this question: graph of parabola going through negative 1, 5 and negative 2, 7. What is the average rate of change from x = −1 to x = −2? 2 −2 4 −1

OpenStudy (anonymous):

HELPPPPp!!!!!!!!!!

OpenStudy (zehanz):

Imagine what happens if you go from x = -1 to x = -2: At x = -1 you are at height 5, at x = -2 you are at height 7.

OpenStudy (anonymous):

2

OpenStudy (zehanz):

True, and because you move only one unit (to the left) the average rate of change would also be 2. But: rate of change is measured going from left to right. The question should have been (imo) what is the average rate of change between x=-2 to x=-1. It is -2.

OpenStudy (anonymous):

That you so much you really are a life saver!!!

OpenStudy (zehanz):

If you like to use a formula: avg rate of change = delta y / delta x, So: \(\dfrac{y_2-y_1}{x_2-x_1}=\dfrac{7-5}{(-2)-(-1)}=\dfrac{2}{-1}=-2\)

OpenStudy (zehanz):

yw!